If the equation for a market demand curve is Qd=15-0.2P and the equation for a market supply curve is Qs=1+ 0.6P, what is the consumer surplus



1
Expert's answer
2021-01-19T12:05:20-0500

A market is in equilibrium when quantity demanded and quantity supplied become equal.

"Q_d = Q_s \\

15 – 0.2P_e = 1 + 0.6P_e \\

14 = 0.8P_e \\

P_e = 17.5 \\

Q_e = 1 + 0.6 \times 17.5 = 11.5"

Consumer surplus is the difference between the maximum price willing to pay and the actual market price. It is the area below the demand curve and above the market price.

Consumer surplus "= 0.5 \times (75 – 17.5) \times 11.5 = 330.6"

Answer: 330.6